a) The first human fo officially run a 4.00 minute mile was DR Roger Bannister in 1954. The current record is 3 minutes 43.13 seconds, set in Rome in 1999 by Hicham El Guerrouj of Morocco. If a milke is 1.60 km, calculate Dr Bannister’s speed in m/s. Show all your steps.
b) A cube is measured to be 32.4 (+)or(-) 0.4 cm each side. Find the volume of the cube including the uncertainty.
Answer #1

A] speed = distance/time = 1mile/4min

= 1.6km/4min

= 1.6*1000 m /[4*60 s] = 6.667 m/s

B] volume = side^3

= 32.4^3 = 34012 cm^3

uncertainty in volume = Volume*3*uncertainty in side/length of side

= 34012*3*0.4/32.4

= 1260 cm^3

Hence volume = 34012 (+/-) 1260 cm^3